OpenStudy (nicia99):

What is the y-intercept of the line given by the equation y equals two x minus nine? The y-intercept is 0. The y-intercept is Negative four point five. The y-intercept is Negative nine. The y-intercept is 18.

OpenStudy (anonymous):

This is an algebra question. Equation:y=2x-9 The constant is your y-intercept. Can you figure it out from that?

OpenStudy (nicia99):

No :(

OpenStudy (anonymous):

To find the y-intercept set x equal to 0. y=2(0)-9 y=0-9 y=-9 (Or you could just look and see that -9 is the constant so it is the y-intercept, but this only works when the equation is equal to y.)

OpenStudy (nicia99):

Ok

OpenStudy (anonymous):

I hope that helped. ☺

OpenStudy (nicia99):

Thanks You Helped Alot :)

OpenStudy (anonymous):

No problem. ☺